# Dynamical scaling study for the estimation of dynamical exponent $z$ of three-dimensional XY spin glass model

### v3

@@ IV Results for 3D ±$\pm J$ XY model
To investigate the critical phenomena of SG systems using the dynamical scaling method, it is necessary to perform a scaling analysis based on Eq. (10); this analysis provides the transition temperature and the product of the critical exponents. For this purpose, identifying a linear size $L$ for which finite-size effects can be neglected within the range of considered MCSs is critical.
@@
First, to examine size dependence, we investigated the behavior near the transition temperature of $\chi_{\mathrm{SG}}$ and $\chi_{\mathrm{CG}}$ defined respectively by Eqs. (6) and (7). These physical quantities exhibit algebraic divergence at the transition temperature. According to Fig. 7, within the range of $5\times 10^{5}$ MCS, the size dependence of $\chi_{\mathrm{SG}}$ disappears for $L\geq 41$ and so does that of $\chi_{\mathrm{CG}}$ for $L\geq 51$. On the basis of these results, the size of the 3D ±$\pm J$ XY model was set to $L=91$ in the present study.
@@
Figure 7: Size dependence of $\chi_{\mathrm{SG}}$ and $\chi_{\mathrm{CG}}$ in the 3D ±$\pm J$ XY model: (a) $\chi_{\mathrm{SG}}$, where size dependence vanishes for $L\geq 41$, and (b) $\chi_{\mathrm{CG}}$, where size dependence vanishes for $L\geq 51$. The temperatures used for the calculations were $T=0.396$ for $\chi_{\mathrm{SG}}$ and $T=0.4460$ for $\chi_{\mathrm{CG}}$, which are close to the critical point.
@@
Second, to estimate the transition temperature, we have to conduct a scaling analysis using the dynamical scaling law represented by Eq. (10). To perform this analysis, it is necessary to determine whether Eq. (13) can be applied to the dynamical scaling of the CG susceptibility. Therefore, we first examine the temperature dependence of $z$ at the CG transition. From the demonstrations described in Sec. III.1 and III.2, we note that the reliability of $z$ obtained from the correlation length estimated by the dynamical scaling law has been established. Based on this result, we calculated the CG correlation function over a temperature range that includes both low- and high-temperature regimes, using $n=256$ and 5 samples. Subsequently, a scaling analysis was conducted according to Eq. (18). In the optimization procedure, short-range correlations ($r<4$ for $f_{\mathrm{CG}}$) and data points with small values ($f_{\mathrm{CG}}(r,t)<2\times 10^{-6}$) were excluded. The initial values of $\{\xi(t)\}$ were set using the correlation length obtained via the second-moment method, as expressed in Eq. (20). The dynamical exponent $z$ was then estimated from the slope of the scaling function. The uncertainties were quantified as the standard deviations of the bootstrap distributions derived from 10,000 resamplings[14]. The result is shown in Fig. 8. This figure clearly shows that, as in the SG case, the CG case also exhibits a dependence of $z$ on $1/T$. Accordingly, Eq. (13) should be employed in the dynamical scaling analysis of the CG susceptibility.
@@
Figure 8: Temperature dependence of $z$ for the CG transition. The dashed line indicates the estimated CG transition temperature obtained in this study.
IV.1 Dynamical Scaling Analysis
@@
Based on the previously obtained temperature dependence of the dynamical exponent $z(T)$, we performed a dynamical scaling analysis using the relation given in Eq. (10).
@@
The quantities used in the calculations for both $\chi_{\mathrm{SG}}$ and $\chi_{\mathrm{CG}}$ were $n=256$, and 5 samples. The temperature range used for the calculations is shown in Fig. 9. In the scaling analysis for both $\chi_{\mathrm{SG}}$ and $\chi_{\mathrm{CG}}$, the dynamical exponent $z$ was treated as temperature dependent, following the assumption described in Eq. (13). As a result, we obtained $T_{\mathrm{SG}}=0.396(1)$ and $T_{\mathrm{CG}}=0.4460(6)$. In addition, the product of critical exponents $\gamma/z\nu$ and $z\nu$ were estimated as scaling parameters. For $\chi_{\mathrm{SG}}$, $\gamma/z\nu=0.3660(4)$ and $z\nu=7.67(5)$ were obtained; for $\chi_{\mathrm{CG}}$, $\gamma/z\nu=0.3214(6)$ and $z\nu=5.70(1)$ were obtained. These results are compared with those reported in previous studies in Table 1 and Table 2.
@@
Although the authors of some previous studies in which the NER method was used did not provide evidence supporting this theory [4, 9], our analysis of the transition temperature using a similar approach yielded results consistent with it. This discrepancy is attributable to finite-time corrections. The computational MCSs in the present study were approximately $5\times 10^{5}$ MCS, which is longer than the MCSs in previous studies[4, 9]. Figure 11 presents the relaxation data obtained in our study, scaled using the parameters from a previous study [9]. The inset shows the distribution of the log-likelihood values obtained via bootstrap sampling, comparing the fits using the parameters from a previous study [9] and those obtained in the present study. To evaluate the log-likelihood functions, we fixed the physical parameters such as $T_{\mathrm{SG}}$, $\gamma/z\nu$, and $z\nu$, and optimized the hyperparameters in the Gaussian process regression for each bootstrap sample. This procedure allows us to evaluate which set of physical parameters is more consistent with the dynamical scaling behavior of the relaxation data. In the prior study, physical quantities were estimated on the basis of relaxation data up to $1\times 10^{5}$ MCS. As shown in Fig. 11, the scaling was successful for data up to $1\times 10^{5}$ MCS. However, when the data were extended to $5\times 10^{5}$ MCS, the scaling failed, which is also clearly reflected in the behavior of the log-likelihood function. This discrepancy is likely attributable to finite-time corrections and supports the validity of our results. Similar issues have been pointed out in finite-size scaling analysis, where the picture can change for small systems[16, 17].
Figure 9: The results of the dynamical scaling analysis for the 3D ±$\pm J$ XY model based on Eq. (10): (a) relaxation data for $\chi_{\mathrm{SG}}$ and (b) the scaling plot. The scaling plot in (b) was obtained using the parameters $T_{\mathrm{SG}}=0.396(1)$, $\gamma/z\nu=0.3660(4)$, and $z\nu=7.67(5)$.
Figure 10: The results of the dynamical scaling analysis for the 3D ±$\pm J$ XY model based on Eq. (10): (a) relaxation data for $\chi_{\mathrm{CG}}$ and (b) the scaling plot. The scaling plot in (b) was generated using the parameters $T_{\mathrm{CG}}=0.4460(6)$, $\gamma/z\nu=0.3214(6)$, and $z\nu=5.70(1)$. In this scaling analysis, $z$ is treated as described in Eq. (13).
@@
IV.2 Estimation of critical exponents
Using the transition temperatures obtained from the analysis in Sec.IV.1, we evaluate the dynamical exponent $z$ at the critical temperatures. Based on these results and the products of critical exponents obtained in Sec.IV.1, we assess the critical behavior.
@@ IV.2 Estimation of critical exponents
From the inference results and Eq. (1), the value of $z$ is evaluated as $z=5.86(1)$ for the SG case and $z=5.24(7)$ for the CG case. The uncertainties were quantified as the standard deviations of the bootstrap distributions derived from 10,000 resamplings[14]. These results and the previously mentioned dynamical scaling analysis in Sec. IV.1 can be used to estimate the critical exponents. These results are compared with those of previous studies in Tables 1 and 2.
The obtained results show improved accuracy for both the transition temperature and the critical exponents compared with previously reported findings. Here, the value of $\gamma/z\nu$ is found to be almost the same as those reported in previous studies employing NER analysis for the SG transition, whereas this is not the case for the CG transition[4, 9]. This discrepancy is likely due to the fact that those earlier studies did not take into account the temperature dependence of the dynamical exponent $z$ in the CG analysis. The dynamical exponent $z$ was found to be lower in the CG case than in the SG case, which supports the result that finite-size effects are more pronounced in the CG case (Fig. 7).
Figure 12: The dynamical SG and CG correlation functions of the 3D ±$\pm J$ XY model at the critical temperature for various Monte Carlo steps $t$: (a) analysis of the SG transition and (b) analysis of the CG transition.
Figure 13: The scaling plots of the SG and CG correlation functions in the 3D ±$\pm J$ XY model: (a) analysis of the SG transition with $\eta_{\mathrm{eff}}=0.6865$, and (b) analysis of the CG transition with $\eta_{\mathrm{eff}}=1.169$.
Figure 14: The SG and CG correlation lengths in the 3D ±$\pm J$ XY model. The dotted line indicates the slope corresponding to the estimated dynamic exponent $z$.
Table 1: Analysis results of the SG transition in the 3D ±$\pm J$ XY model compared with the corresponding results reported in previous studies.
@@
Present work 0.396(1) 2.80(3) 5.86(1) 1.30(4)
Table 2: Analysis results of the CG transition in the 3D ±$\pm J$ XY model compared with those in previous studies.
@@
Present work 0.4460(6) 2.54(1) 5.24(7) 1.08(4)
V Summary and Discussion
We primarily aimed to validate a method for calculating the dynamical exponent from the correlation function, thereby enabling an analysis of SG systems. The obtained dynamical exponent appeared to be reliable, showed good accuracy, and was not limited to SG systems. The results confirm the reliability of this method. Using this approach, we analyzed the 3D ±$\pm J$ XY model. Compared with conventional methods, our method achieved a higher precision in evaluating the dynamical exponent. Through this analysis, we also examined the temperature dependence of $z(T)$ in the CG transition. In addition, our results support the validity of the spin chirality decoupling picture in SG systems.
@@ V Summary and Discussion
The correlations estimated using our method also face problems related to finite-time corrections. Because we do not extrapolate the scaling regime as $t\rightarrow\infty$ for the calculation of the dynamical exponent, our proposed method would suffer a difficult to apply to cases with strong fluctuations. An example of such a situation is the analysis of SG transitions near multi-critical points in the 3D ±$\pm J$ Ising model[2]. Making the method more widely applicable will require the development of an extrapolation method for $t\rightarrow\infty$. A reliable and systematic extrapolation scheme has recently been proposed in the NER method[19]. We expect that using such a scheme will lead to solutions to such problems. This work is also a future task.

### v4

@@ 4 Results for 3D $\pm J$ XY model
To investigate the critical phenomena of SG systems using the dynamical scaling method, it is necessary to perform a scaling analysis based on Eqs. (10) and (14); this analysis provides the transition temperature and the product of the critical exponents. For this purpose, identifying a linear size $L$ for which finite-size effects can be neglected within the range of considered MCSs is critical.
@@
First, to examine size dependence, we investigated the behavior near the transition temperature of $\chi_{\mathrm{SG}}$ and $\chi_{\mathrm{CG}}$ defined respectively by Eqs. (6) and (7). These physical quantities exhibit algebraic divergence at the transition temperature. According to Fig. 8, within the range of $5\times 10^{5}$ MCS, the size dependence of $\chi_{\mathrm{SG}}$ disappears for $L\geq 41$ and so does that of $\chi_{\mathrm{CG}}$ for $L\geq 51$. On the basis of these results, the size of the 3D $\pm J$ XY model was set to $L=91$ in the present study.
@@
Figure 8: Size dependence of $\chi_{\mathrm{SG}}$ and $\chi_{\mathrm{CG}}$ in the 3D $\pm J$ XY model: (a) $\chi_{\mathrm{SG}}$, where size dependence vanishes for $L\geq 41$, and (b) $\chi_{\mathrm{CG}}$, where size dependence vanishes for $L\geq 51$. The temperatures used for the calculations were $T=0.379$ for $\chi_{\mathrm{SG}}$ and $T=0.4460$ for $\chi_{\mathrm{CG}}$, which are close to the critical point.
Second, to estimate the transition temperature, we have to conduct a scaling analysis using the dynamical scaling law represented by Eqs. (10) and (14).
@@
To perform this analysis, it is necessary to examine the temperature dependence of $z$ expressed in Eqs. 13 and 17. From the demonstrations described in Secs. 3.1 and 3.2, we note that the reliability of $z$ obtained from the correlation length estimated by the dynamical scaling law has been established. Based on this result, we calculated the SG and CG correlation functions over a temperature range that includes both low- and high-temperature regimes, using $n=256$ and 5 samples. Subsequently, a scaling analysis was conducted according to Eq. (22).
In an optimization procedure, we discarded short-range correlations ($r<3$ for $f_{\mathrm{SG}}$ and $r<4$ for $f_{\mathrm{CG}}$) and data of small values ($f_{\mathrm{SG}}(r,t)<10^{-5}$ and $f_{\mathrm{CG}}(r,t)<2\times 10^{-6}$).
The dynamical exponent $z$ was then estimated from the slope of the scaling function. The uncertainties were quantified as the standard deviations of the bootstrap distributions derived from 128 resamplings [15].
The result is shown in Fig. 9. According to this figure, the temperature dependence of $z$ expressed in Eqs. 13 and 17 is confirmed for both SG and CG susceptibilities. Although it has been reported in other SG systems that this linear temperature dependence breaks down at low temperatures [7], it is considered to remain consistent in the vicinity of the transition temperature. In addition, the intercepts obtained from linear fits to the data do not necessarily vanish; however, the linear fits constructed from the transition temperature and the value of $z$ at the transition point, with the intercept fixed to zero, provide good approximations. These observations indicate that the dynamical scaling analysis based on Eqs. 13 and 17 is required.
@@
Figure 9: Temperature dependence of $z$ for the SG and CG transitions. (a) shows the result for the SG case, and (b) shows the result for the CG case. The dashed line indicates the estimated transition temperatures obtained in this study. The linear fits of $z$ are plotted using the transition temperatures and the values of $z$ at the corresponding critical points.
4.1 Dynamical Scaling Analysis
@@
Based on the previously obtained temperature dependence of the dynamical exponent $z(T)$, we performed a dynamical scaling analysis using the relation given in Eqs. (10) and (14).
The quantities used in the calculations for both $\chi_{\mathrm{SG}}$ and $\chi_{\mathrm{CG}}$ were $n=256$ and 5 samples. The temperature range used for the calculations is shown in Figs. 10 and 11. For $\chi_{\mathrm{SG}}$, the temperature range was chosen such that the dynamical scaling described by Eq. (10) does not break down and remains valid in the vicinity of the critical temperature. Specifically, including data down to $T\geq 0.410$ yields an apparent critical temperature of $T=0.4143(2)$ in the dynamical scaling analysis, which contradicts the fact that Eqs. (10) are applicable only on the high-temperature side. To avoid this inconsistency while still covering temperatures sufficiently close to the transition, we adopt the range $T\geq 0.420$.
@@
In the scaling analysis for both $\chi_{\mathrm{SG}}$ and $\chi_{\mathrm{CG}}$, the dynamical exponent $z$ was treated as temperature dependent, following the assumption described in Eqs. (13) and (17). As a result, we obtained $T_{\mathrm{SG}}=0.379(3)$ and $T_{\mathrm{CG}}=0.4460(6)$. In addition, the product of critical exponents $\gamma/z\nu$ and $z\nu$ were estimated as scaling parameters. For $\chi_{\mathrm{SG}}$, $\gamma/z\nu=0.3513(13)$ and $z\nu=8.65(14)$ were obtained; for $\chi_{\mathrm{CG}}$, $\gamma/z\nu=0.3214(6)$ and $z\nu=5.70(1)$ were obtained. These results are compared with those reported in previous studies in Table 1 and Table 2.
@@
Although the authors of some previous studies in which the NER method was used did not provide evidence supporting this theory [4, 9], our analysis of the transition temperature using a similar approach yielded results consistent with it. This discrepancy is attributable to finite-time corrections. The computational MCSs in the present study were approximately $5\times 10^{5}$ MCS, which is longer than the MCSs in previous studies[4, 9]. Figure 12 presents the relaxation data obtained in our study, scaled using the parameters from a previous study [9]. Since the temperature used in the previous study was $T=0.435$, we adopted the temperature range $T\geq 0.460$ for this comparison. The inset shows the distribution of the log-likelihood values obtained via bootstrap sampling, comparing the fits using the parameters from a previous study [9] and those obtained in the present study. To evaluate the log-likelihood functions, we fixed the physical parameters such as $T_{\mathrm{SG}}$, $\gamma/z\nu$, and $z\nu$, and optimized the hyperparameters in the Gaussian process regression for each bootstrap sample. This procedure allows us to evaluate which set of physical parameters is more consistent with the dynamical scaling behavior of the relaxation data. In the prior study, physical quantities were estimated on the basis of relaxation data up to $1\times 10^{5}$ MCS. As shown in Fig. 12, the scaling was successful for data up to $1\times 10^{5}$ MCS. However, when the data were extended to $5\times 10^{5}$ MCS, the scaling failed, which is also clearly reflected in the behavior of the log-likelihood function. This discrepancy is likely attributable to finite-time corrections and supports the validity of our results. Similar issues have been pointed out in finite-size scaling analysis, where the picture can change for small systems[17, 18].
Figure 10: The results of the dynamical scaling analysis for the 3D $\pm J$ XY model based on Eq. (10): (a) relaxation data for $\chi_{\mathrm{SG}}$ and (b) the scaling plot. The scaling plot in (b) was obtained using the parameters $T_{\mathrm{SG}}=0.379(3)$, $\gamma/z\nu=0.3513(13)$, and $z\nu=8.65(14)$.
Figure 11: The results of the dynamical scaling analysis for the 3D $\pm J$ XY model based on Eq. (14): (a) relaxation data for $\chi_{\mathrm{CG}}$ and (b) the scaling plot. The scaling plot in (b) was generated using the parameters $T_{\mathrm{CG}}=0.4460(6)$, $\gamma/z\nu=0.3214(6)$, and $z\nu=5.70(1)$. In this scaling analysis, $z$ is treated as described in Eq. (17).
@@
4.2 Estimation of critical exponents
Using the transition temperatures obtained from the analysis in Sec.4.1, we evaluate the dynamical exponent $z$ at the critical temperatures. Based on these results and the products of critical exponents obtained in Sec.4.1, we assess the critical behavior.
@@ 4.2 Estimation of critical exponents
From the inference results and Eq. (1), the value of $z$ is evaluated as $z=6.148(1)$ for the SG case and $z=5.145(4)$ for the CG case. The uncertainties were quantified as the standard deviations of the bootstrap distributions derived from 100 resamplings[15]. These results and the previously mentioned dynamical scaling analysis in Sec. 4.1 can be used to estimate the critical exponents. These results are compared with those of previous studies in Tables 1 and 2.
The obtained results show improved accuracy for both the transition temperature and the critical exponents compared with previously reported findings.
Here, the value of $\gamma/z\nu$ is found to be almost the same as those reported in previous studies employing NER analysis for the SG and CG transition.
From Fig. 8, one can also extract the value of $\gamma/z\nu$ from the slope of $\chi_{\mathrm{SG}}$ at the transition temperature. Using this procedure, we find $\gamma/z\nu=0.315$ for the SG case and $\gamma/z\nu=0.313$ for the CG case. This demonstrates that the deviation from the value obtained by the dynamical scaling analysis is larger in the SG case than in the CG case. These results suggest that the finite-time corrections affect the dynamical-scaling analysis and the susceptibility analysis to different degrees, indicating the necessity of an extrapolation toward the $t\to\infty$ limit in order to obtain more reliable estimates.
In addition, the dynamical exponent $z$ was found to be lower in the CG case than in the SG case, which supports the observation that finite-size effects are more pronounced in the CG sector (Fig. 8).
Figure 13: The dynamical SG and CG correlation functions of the 3D $\pm J$ XY model at the critical temperature for various Monte Carlo steps $t$: (a) analysis of the SG transition and (b) analysis of the CG transition.
Figure 14: The scaling plots of the SG and CG correlation functions in the 3D $\pm J$ XY model: (a) analysis of the SG transition with $\eta_{\mathrm{eff}}=0.210(3)$, and (b) analysis of the CG transition with $\eta_{\mathrm{eff}}=0.510(3)$.
Figure 15: The SG and CG correlation lengths in the 3D $\pm J$ XY model. The dotted line indicates the slope corresponding to the estimated dynamic exponent $z$.
Table 1: Analysis results of the SG transition in the 3D $\pm J$ XY model compared with the corresponding results reported in previous studies.
@@
Present work 0.379(3) 3.03(5) 6.148(1) 1.40(2)
Table 2: Analysis results of the CG transition in the 3D $\pm J$ XY model compared with those in previous studies.
@@
Present work 0.4460(6) 1.831(5) 5.145(4) 1.10(1)
5 Summary and Discussion
We primarily aimed to validate a method for calculating the dynamical exponent from the correlation function, thereby enabling an analysis of SG systems. The obtained dynamical exponent appeared to be reliable, showed good accuracy, and was not limited to SG systems. The results confirm the reliability of this method. Using this approach, we analyzed the 3D $\pm J$ XY model. Compared with conventional methods, our method achieved a higher precision in evaluating the dynamical exponent. Through this analysis, we also examined the temperature dependence of $z(T)$ in the CG transition. In addition, our results support the validity of the spin chirality decoupling picture in SG systems.
@@ 5 Summary and Discussion
The correlations estimated using our method also face problems related to finite-time corrections. Because we do not extrapolate the scaling regime as $t\rightarrow\infty$ for the calculation of the dynamical exponent, our proposed method would suffer a difficult to apply to cases with strong fluctuations. An example of such a situation is the analysis of SG transitions near multi-critical points in the 3D $\pm J$ Ising model[2]. Making the method more widely applicable will require the development of an extrapolation method for $t\rightarrow\infty$. A reliable and systematic extrapolation scheme has recently been proposed in the NER method[16]. We expect that using such a scheme will lead to solutions to such problems. This work is also a future task.